# Categorical Reparameterization with Denoising Diffusion models

### v1

@@ 3 Method
In this section we present ReDGE (Reparameterized Diffusion Gradient Estimator), which builds upon diffusion models to define an approximate pathwise gradient for $\pi_{\!\theta}$. We start by recalling the basics of these models.
@@
We present denoising diffusion models (DDMs) (sohl2015deep; song2019generative; ho2020denoising) and the DDIM framework (song2021ddim) through the interpolation viewpoint (liu2023flow; lipman2023flow; albergo2023stochastic). We provide more details in Appendix C. DDMs define a generative procedure for a data distribution $\pi_{0}$ by first specifying a continuous family of marginals $(\pi_{t})_{t\in[0,1]}$ that connects $\pi_{0}$ to the simple reference distribution $\pi_{1}\coloneqq\mathcal{N}(0,\mathbf{I})$. More precisely, $\pi_{t}=\text{Law}(X_{t})$, where
$X_{t}=\alpha_{t}X_{0}+\sigma_{t}X_{1}\;,\quad X_{0}\sim\pi_{0}\;,\quad X_{1}\sim\pi_{1}\;.$ (8)
Here $X_{0}$ and $X_{1}$ are independent and $(\alpha_{t})_{t\in[0,1]}$ and $(\sigma_{t})_{t\in[0,1]}$ are non-increasing and non-decreasing, respectively, schedules with boundary conditions $(\alpha_{0},\sigma_{0})\coloneqq(1,0)$ and $(\alpha_{1},\sigma_{1})\coloneqq(0,1)$. A popular example is the linear schedule, defined by $(\alpha_{t},\sigma_{t})=(1-t,t)$ (lipman2023flow; esser2024scaling). To generate new samples, DDMs simulate a time-reversed Markov chain. Given a decreasing sequence $({t_{k}})_{k=0}^{n-1}$ of $n$ time steps with ${t_{n-1}}=1$ and ${t_{0}}=0$, reverse transitions are iteratively applied to map a sample from $\pi_{{t_{k+1}}}$ to one from $\pi_{{t_{k}}}$, thereby progressively denoising until the clean data distribution $\pi_{0}$ is reached.
The DDIM framework (song2021ddim) introduces a general family of reverse transitions for denoising diffusion models. It relies on a schedule $(\eta_{t})_{t\in[0,1]}$, satisfying $\eta_{t}\leq\sigma_{t}$ for all $t\in[0,1]$, along with a family of conditional distribution given for $s<t$ by
@@
When $\eta_{s}=0$, this Gaussian is understood, by abuse of notation, as a Dirac delta centered at the same mean. Clearly, for all $\eta_{s}\in[0,\sigma_{s}]$, a sample from $q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\cdot|X_{0},X_{1})$ with $(X_{0},X_{1})\sim\pi_{0}\otimes\mathcal{N}(0,\mathbf{I})$ is a sample from $\pi_{s}$. We define the reverse transition
$\pi^{\eta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{s}|x_{t})$ $=\mathbb{E}\Big[\,q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(x_{s}|X_{0},X_{1})\,\Big|\,X_{t}=x_{t}\Big]$ (9)
$=\mathbb{E}\Big[\,q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(x_{s}|X_{0},\frac{x_{t}-\alpha_{t}X_{0}}{\sigma_{t}})\,\Big|\,X_{t}=x_{t}\Big]$
where the joint distribution of the random variables $(X_{0},X_{t},X_{1})$ is defined in (8) and in the second line we have used that the $X_{1}|X_{0},X_{t}\sim\updelta_{(X_{t}-\alpha_{t}X_{0})/\sigma_{t}}$. For simplicity, we define $q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,t}}(x_{s}|x_{0},x_{t})\coloneqq q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(x_{s}|x_{0},(x_{t}-\alpha_{t}x_{0})/\sigma_{t})$. By construction, the transitions (9) satisfy the marginalization property $\smash{\pi_{s}(x_{s})=\int\pi^{\eta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{s}|x_{t})\,\pi_{t}(x_{t})\mathrm{d}x_{t}}$. Thus, $(\pi^{\eta}_{{t_{k}}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{k+1}}})_{k=0}^{n-2}$ defines a set of reverse transitions that enable stepwise sampling from the sequence $(\pi_{t_{k}})_{k=0}^{n-1}$. In practice, however, these transitions are intractable. A common approximation is to replace $X_{0}$ in the second line of (9) by its conditional expectations (ho2020denoising; song2021ddim). More precisely, let $\hat{x}_{0}(x_{t},t)\coloneqq\int x_{0}\,\pi_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{0}|x_{t})\mathrm{d}x_{0}$, where $\pi_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}$ is defined as the conditional distribution of $X_{0}$ given $X_{t}$ in (8). Then the model proposed in (ho2020denoising; song2021ddim) corresponds to approximating each $\pi^{\eta}_{t_{k}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t_{k+1}}$ by
$\hat{\pi}^{\eta}_{k\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}k+1}(x_{k}|x_{k+1})\coloneqq q^{\eta}_{\smash{{t_{k}}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,{t_{k+1}}}}(x_{k}|\hat{x}_{0}(x_{k+1},{t_{k+1}}),x_{k+1}).$
@@ 3.2 Diffusion-based categorical reparameterization
1:Input: grid $({t_{k}})_{k=0}^{n-1}$, schedules $(\alpha_{{t_{k}}},\sigma_{{t_{k}}},\eta_{{t_{k}}})_{k=0}^{n-1}$
@@
6:  $\mu\leftarrow\alpha_{{t_{k}}}\hat{x}_{0}+(\sigma_{{t_{k}}}^{2}-\eta_{{t_{k}}}^{2})^{1/2}\hat{x}_{1}$
7:  $x\leftarrow\mu+\eta_{{t_{k}}}z$   with $z\sim\mathcal{N}(0,\mathbf{I}_{K})^{\otimes L}$
8:end for
9:return $x$
We now introduce our diffusion-based soft reparameterization of $\pi_{\!\theta}$. This reparameterization is built upon a DDM with target $\pi^{\theta}_{0}=\pi_{\!\theta}$. Since $\pi_{\!\theta}$ is a discrete measure, the resulting denoising distribution denoted by $\pi^{\theta}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}$ is itself discrete. Indeed, following (8) and the factorization (2), $\pi^{\theta}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{0}|x_{t})\propto\prod_{i=1}^{L}\pi^{\theta,i}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x^{i}_{0}|x^{i}_{t})$ where
@@
With this structure, the posterior-mean denoiser $\hat{x}^{\theta}_{0}(x_{t},t)\coloneqq\sum_{x_{0}}x_{0}\,\pi^{\theta}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{0}|x_{t})$ simplifies to a matrix of posterior probabilities because of the one-hot structure; i.e. we have for any $i\in[L]$ and $j\in[K]$ that $\hat{x}^{\theta}_{0}(x_{t},t)^{ij}=\pi^{\theta,i}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(e_{j}|x_{t})$ and the denoiser can be computed exactly and efficiently. Indeed, since $\|x^{i}_{t}-\alpha_{t}e_{j}\|^{2}=\|x^{i}_{t}\|^{2}-2\alpha_{t}x^{ij}_{t}+\alpha^{2}_{t}$, we get
@@
$=\frac{\pi^{i}_{\!\theta}(e_{j})\exp(\alpha_{t}x^{ij}_{t}/\sigma^{2}_{t})}{\sum_{k=1}^{K}\pi^{i}_{\!\theta}(e_{k})\exp(\alpha_{t}x^{ik}_{t}/\sigma^{2}_{t}))}\;.$
This yields the following simple matrix form for the denoiser: $\hat{x}^{\theta}_{0}(x_{t},t)=\text{softmax}(\varphi_{\theta}+\alpha_{t}x_{t}/\sigma^{2}_{t})\;.\vskip 2.84544pt$ (10) Unlike standard diffusion models that learn an approximate denoiser via a neural network, here the denoiser $\hat{x}^{\theta}_{0}(\cdot,t)$ admits a closed-form expression thanks to the factorized categorical structure. This allows reverse transitions from $\pi_{1}$ to $\pi_{\!\theta}$ free of denoiser approximation and results in an approximate and differentiable sampling procedure. For simplicity we consider only the deterministic sampler corresponding to $\eta_{s}=0$ for all $s\in[0,1]$. Define
$T^{\theta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{t})\coloneqq(\alpha_{s}-\alpha_{t}\sigma_{s}/\sigma_{t})\hat{x}^{\theta}_{0}(x_{t},t)+{\sigma_{s}}x_{t}/\sigma_{t}\;.$ (11)
$\smash{T^{\theta}_{t_{k}}(x_{1})\coloneqq T^{\theta}_{t_{k}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t_{k+1}}\circ\dotsc\circ T^{\theta}_{t_{n-2}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t_{n-1}}(x_{1})}\;.$ (12)
Then, $T^{\theta}_{t_{k}}(X_{1})$ with $X_{1}\sim\mathcal{N}(0,\mathbf{I}_{K})^{\otimes L}$ is an approximate sample from the Gaussian mixture with density $\pi^{\theta}_{\!\,{t_{k}}}(x_{t_{k}})\coloneqq\sum_{x_{0}}\prod_{i=1}^{L}\mathrm{N}(x^{i}_{t_{k}};\alpha_{t_{k}}x^{i}_{0},\sigma^{2}_{t_{k}}\mathbf{I}_{K})\pi_{\!\theta}(x_{0})$ and in particular, $T^{\theta}_{0}(X_{1})$ is the approximate and relaxed sample from $\pi_{\!\theta}$ that we use to compute the gradient estimator in (7).
Note that with a single diffusion step, the reparameterized sample is $T^{\theta}_{0}(X_{1})=\hat{x}^{\theta}_{0}(X_{1},1)=\mathbb{E}_{\pi_{\!\theta}}[X_{0}]$, because of the boundary condition $\alpha_{1}=0$, and we recover the Straight-Through estimator. On the other hand, using many diffusion steps and well-placed timesteps $({t_{k}})_{k=0}^{n-1}$ yields an almost exact reparameterization of $\pi_{\!\theta}$. As discussed previously, this is precisely what we want to avoid: the mapping becomes nearly piecewise constant in $\theta$ and we end up with a high-variance reparameterized gradient. This trade-off is directly analogous to the temperature parameter $\tau$ in Gumbel-Softmax relaxations, where a high temperature yields a relaxed but biased approximation whereas a low temperature recovers a high-variance estimator. In our case, the role of the relaxation parameter is played by the number of diffusion steps and the placement of the timesteps $({t_{k}})_{k=1}^{n-1}$.
More precisely, we show in Proposition 3.1 that the earliest timestep $t_{1}$} governs the behaviour of the reparameterized gradient: as ${t_{1}}\to 0$, the last DDIM step $T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}$ collapses almost all points in ${\mathbb{R}}^{K}$ onto a single one–hot vector, and as consequence, the Jacobian of $T^{\theta}_{0}$ w.r.t. $\theta$ vanishes. We make this intuition precise by considering next, and w.l.o.g, the case $L=1$ and $\varphi_{\theta}=\theta\in{\mathbb{R}}^{K}$. We consider the DDIM map $T^{\theta}_{0}$ as a function of the input noise $x_{1}$ and ${t_{1}}\in(0,1)$ while the remaining steps $t_{k}$} with $k\geq 2$ are fixed. For this reason we make the dependence on the timesteps $t_{1:n-1}$} explicit and write $(x_{1},{t_{1:n-1}})\mapsto T^{\theta}_{0}(x_{1};{t_{1:n-1}})$ for $T^{\theta}_{0}$. The proof is given in Appendix A.
Proposition 3.1 (Informal).
@@
Under assumptions stated in the Appendix, and with the timesteps $({t_{k}})_{k=2}^{n-1}$ fixed, we have for all $\theta\in\Theta$, $\lim_{{t_{1}}\to 0}\big\|\operatorname{J}_{\theta}\!T^{\theta}_{0}(X_{1};{t_{1:n-1}})\big\|=0\;,\quad\mathbb{P}-\text{a.s.}$ (13) with $X_{1}\sim\mathcal{N}(0,\mathbf{I}_{K})$. The proof relies on the fact that for any $t\in(0,1]$ and $x\in{\mathbb{R}}^{K}$
$\begin{cases}\operatorname{J}_{\theta}\!T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x)&=\Sigma^{\theta}_{t}(x),\\ \operatorname{J}_{x}\!T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x)&=\alpha_{t}\Sigma^{\theta}_{t}(x)\big/\sigma^{2}_{t}\;,\end{cases}$ (14)
with $\Sigma^{\theta}_{t}(x)\!\coloneqq\!\mathbb{C}\mathrm{ov}_{\pi^{\theta}_{0|t}(\cdot|x)}(X_{0})$. $\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(\cdot|x)$ is a categorical distribution with probability vector $\text{softmax}(\theta+\alpha_{t}x/\sigma^{2}_{t})$ and under the assumption that $\alpha_{t}/\sigma^{2}_{t}\rightarrow\infty$ as $t\rightarrow 0$ it collapses into a Dirac delta unless $x$ has at least two coordinates equal to $\max_{i}x^{i}$. This leads us to consider the decision boundary $\mathsf{H}:=\bigcup_{j\neq k}\{x\in\mathbb{R}^{K}:x^{j}=x^{k}=\max_{i}x^{i}\}.$ Outside of this set, the norm of $\Sigma^{\theta}_{t}(x)$ goes to 0 when $t\to 0$ as fast as $\exp(-\alpha_{t}L(x,\theta)/\sigma^{2}_{t})$ where $L(x,\theta)>0$. Thus, both the Jacobians (14) go to 0. We then get the result by assuming that the limit, as ${t_{1}}\to 0$, of the DDIM trajectories $T^{\theta}_{{t_{1}}}(X_{1})$, with $X_{1}\sim\mathcal{N}(0,\mathbf{I}_{K})$, land outside of $\mathsf{H}$ almost surely. We illustrate Proposition 3.1 in Figure 1.
@@
Following the previous discussion, the timestep $t_{1}$} must be therefore be chosen in an intermediate regime, small enough to reduce the bias but not so small that the gradients become uninformative.
3.3 Hard gradient estimator.
A natural choice of hard gradient estimator is
$\operatorname{J}_{\theta}\!T^{\theta}_{0}(X_{1})^{\top}\nabla_{\!{x}}f(X_{0})$ (15)
where $\smash{X_{0}\sim\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(\cdot|T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(X_{1}))}$; i.e. we draw a hard sample $X_{0}$ only at the last diffusion step.

### v2

@@ 3 Method
In this section, we present ReDGE (Reparameterized Diffusion Gradient Estimator), which builds on diffusion models to define a continuous relaxation for $\pi_{\!\theta}$. We begin by reviewing the basics of these models.
@@
We present denoising diffusion models (DDMs) (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020) and the DDIM framework (Song et al., 2021) from the interpolation viewpoint (Liu et al., 2023b; Lipman et al., 2023; Albergo et al., 2023). More details are provided in Appendix C.
DDMs define a generative procedure for a data distribution $\pi_{0}$ by specifying a continuous family of marginals $(\pi_{t})_{t\in[0,1]}$ that connects $\pi_{0}$ to the simple reference distribution $\pi_{1}\coloneqq\mathcal{N}(0,\mathbf{I})$. More precisely, we consider here $\pi_{t}=\text{Law}(X_{t})$, where
$X_{t}=\alpha_{t}X_{0}+\sigma_{t}X_{1}\;,$ (7)
$X_{0}$ and $X_{1}$ are independent samples from $\pi_{0}$ and $\pi_{1}$ respectively. In addition, $(\alpha_{t})_{t\in[0,1]}$ and $(\sigma_{t})_{t\in[0,1]}$ are non-increasing and non-decreasing schedules, respectively, with boundary conditions $(\alpha_{0},\sigma_{0})\coloneqq(1,0)$ and $(\alpha_{1},\sigma_{1})\coloneqq(0,1)$. To generate new samples, DDMs simulate a time-reversed Markov chain. Given a decreasing sequence $({t_{k}})_{k=0}^{n-1}$ of $n$ time steps with ${t_{n-1}}=1$ and ${t_{0}}=0$, reverse transitions are applied iteratively to map a sample from $\pi_{{t_{k+1}}}$ to one from $\pi_{{t_{k}}}$, progressively denoising until the clean data distribution $\pi_{0}$ is reached.
The DDIM framework (Song et al., 2021) introduces a general family of reverse transitions for denoising diffusion models. It relies on a schedule $(\eta_{t})_{t\in[0,1]}$, satisfying $\eta_{t}\leq\sigma_{t}$ for all $t\in[0,1]$, along with a family of conditional distribution given for $s<t$ by
@@
When $\eta_{s}=0$, this Gaussian is understood, by abuse of notation, as a Dirac mass centered at the same mean. Clearly, for all $\eta_{s}\in[0,\sigma_{s}]$, a sample from $q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\cdot|X_{0},X_{1})$ with $(X_{0},X_{1})\sim\pi_{0}\otimes\mathcal{N}(0,\mathbf{I})$ is a sample from $\pi_{s}$. Note that if $X_{s}^{\eta}|X_{0},X_{1}\sim q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\cdot|X_{0},X_{1})$, then $X_{s}^{\eta}|X_{0},X_{t}\sim q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,t}}(\cdot|x_{0},x_{t})=q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,1}}(\cdot|X_{0},(X_{t}-\alpha_{t}X_{0})/\sigma_{t})$ where the joint distribution of the random variables $(X_{0},X_{t},X_{1})$ is defined in (7). We define the reverse transition
$\pi^{\eta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{s}|x_{t})$ $=\mathbb{E}\Big[\,q^{\eta}_{\smash{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,t}}(x_{s}|X_{0},X_{t})\,\Big|\,X_{t}=x_{t}\Big]\;.$ (8)
By construction, the transitions (8) satisfy the marginalization property, i.e. for any $0\leq s<t\leq 1$, $\smash{\pi_{s}(x_{s})=\int\pi^{\eta}_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{s}|x_{t})\,\pi_{t}(x_{t})\mathrm{d}x_{t}}$. Thus, $(\pi^{\eta}_{{t_{k}}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{k+1}}})_{k=0}^{n-2}$ defines a set of reverse transitions that enable stepwise sampling from the sequence $(\pi_{t_{k}})_{k=0}^{n-1}$. In practice, however, these transitions are intractable. A common approximation is to replace $X_{0}$ in the second line of (8) by its conditional expectations (Ho et al., 2020; Song et al., 2021). More precisely, let $\hat{x}_{0}(x_{t},t)\coloneqq\int x_{0}\,\pi_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{0}|x_{t})\mathrm{d}x_{0}$, where $\pi_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}$ is defined as the conditional distribution of $X_{0}$ given $X_{t}$ in (7). Then the model proposed in (Ho et al., 2020; Song et al., 2021) corresponds to approximating each $\pi^{\eta}_{t_{k}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t_{k+1}}$ by
$\hat{p}^{\eta}_{k\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}k+1}(x_{k}|x_{k+1})\coloneqq q^{\eta}_{\smash{{t_{k}}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}0,{t_{k+1}}}}(x_{k}|\hat{x}_{0}(x_{k+1},{t_{k+1}}),x_{k+1}).$
For simplicity, we consider next only the deterministic sampler with $\eta_{s}=0$ for all $s\in[0,1]$. Then $\hat{p}^{\eta}_{k\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}k+1}(\cdot|x_{k+1})$ becomes a Dirac at $T_{t_{k}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t_{k+1}}(x_{t_{k+1}})$ where for $s<t$:
$T_{s\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{t})\coloneqq(\alpha_{s}-\alpha_{t}\sigma_{s}/\sigma_{t})\hat{x}_{0}(x_{t},t)+{\sigma_{s}}x_{t}/\sigma_{t}\;.$ (9)
$\smash{T_{t_{k}}(x_{1})\coloneqq T_{t_{k}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t_{k+1}}\circ\dotsc\circ T_{t_{n-2}\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t_{n-1}}(x_{1})}\;.$ (10)
@@ 3.2 Diffusion-based categorical reparameterization
1:Input: grid $({t_{k}})_{k=0}^{n-1}$, schedule $(\alpha_{{t_{k}}},\sigma_{{t_{k}}})_{k=0}^{n-1}$
@@
6:  $x\leftarrow\alpha_{{t_{k}}}\hat{x}_{0}+\sigma_{{t_{k}}}\hat{x}_{1}$
7:end for
8:return $x$
We now introduce our diffusion-based soft reparameterization of $\pi_{\!\theta}$. This reparameterization is based on a DDM with target $\pi^{\theta}_{0}=\pi_{\!\theta}$. Since $\pi_{\!\theta}$ is a discrete measure, the resulting denoising distribution, denoted by $\pi^{\theta}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}$, is also discrete. Indeed, by (7) and the factorization (2), the conditional distribution factorizes as $\pi^{\theta}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{0}|x_{t})\propto\prod_{i=1}^{L}\pi^{\theta,i}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x^{i}_{0}|x^{i}_{t})$, where
@@
With this structure, the posterior-mean denoiser $\hat{x}^{\theta}_{0}(x_{t},t)\coloneqq\sum_{x_{0}}x_{0}\,\pi^{\theta}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(x_{0}|x_{t})$ simplifies to a matrix of posterior probabilities due to the one-hot structure; that is, for any $i\in[L]$ and $j\in[K]$, we have $\hat{x}^{\theta}_{0}(x_{t},t)^{ij}=\pi^{\theta,i}_{\!\,0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}t}(e_{j}|x_{t})$, and the denoiser can be computed exactly and efficiently. Indeed, since $\|x^{i}_{t}-\alpha_{t}e_{j}\|^{2}=\|x^{i}_{t}\|^{2}-2\alpha_{t}x^{ij}_{t}+\alpha^{2}_{t}$, we get
@@
$=\frac{\exp(\varphi_{\theta}^{ij})\exp(\alpha_{t}x^{ij}_{t}/\sigma^{2}_{t})}{\sum_{k=1}^{K}\exp(\varphi_{\theta}^{ik})\exp(\alpha_{t}x^{ik}_{t}/\sigma^{2}_{t}))}\;.$
This yields the following simple matrix form for the denoiser: $\hat{x}^{\theta}_{0}(x_{t},t)=\text{softmax}(\varphi_{\theta}+\alpha_{t}x_{t}/\sigma^{2}_{t})\;.\vskip 2.84544pt$ (11) Unlike standard diffusion models that learn an approximate denoiser using a neural network, here the denoiser $\hat{x}^{\theta}_{0}(\cdot,t)$ has a closed-form expression due to the factorized categorical structure. This enables reverse transitions from $\pi_{1}$ to $\pi_{\!\theta}$ without denoiser approximation and yields an approximate, differentiable sampling procedure. Denote for any k<n2k<n-2 by $T^{\theta}_{t_{k}}$ the DDIM map associated with $\hat{x}^{\theta}_{0}(x_{t},t)$ defined in (9). Then, $T^{\theta}_{t_{k}}(X_{1})$ with $X_{1}\sim\mathcal{N}(0,\mathbf{I}_{K})^{\otimes L}$ is an approximate sample from the Gaussian mixture with density $\pi^{\theta}_{\!\,{t_{k}}}(x_{t_{k}})\coloneqq\sum_{x_{0}}\prod_{i=1}^{L}\mathrm{N}(x^{i}_{t_{k}};\alpha_{t_{k}}x^{i}_{0},\sigma^{2}_{t_{k}}\mathbf{I}_{K})\pi_{\!\theta}(x_{0})$ and $T^{\theta}_{0}(X_{1})$ is an approximate relaxed sample from $\pi_{\!\theta}$. By (6), a natural choice of gradient estimator is $\operatorname{J}_{\theta}\!T^{\theta}_{0}(X_{1})^{\top}\nabla_{\!{x}}f(X_{0})\;,\quad$ (ReDGE) where $\smash{X_{0}\sim\pi^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(\cdot|T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}(X_{1}))}$. As a result, with a single diffusion step, the reparameterized sample is $T^{\theta}_{0}(X_{1})=\hat{x}^{\theta}_{0}(X_{1},1)=\mathbb{E}_{\pi_{\!\theta}}[X_{0}]$, due to the boundary condition $\alpha_{1}=0$, and we recover the Straight-Through estimator (both soft and hard) as a special case. In contrast, using many diffusion steps and appropriately chosen timesteps $({t_{k}})_{k=0}^{n-1}$ yields an almost exact reparameterization of $\pi_{\!\theta}$. As discussed previously, this is precisely the regime we seek to avoid: the gradient of the mapping essentially vanishes, resulting in a high-variance reparameterized gradient. This trade-off is directly analogous to the role of the temperature parameter $\tau$ in Gumbel-Softmax relaxations, where a high temperature yields a relaxed but biased approximation, while a low temperature results in a high-variance estimator. In our case, the relaxation parameter is determined by the number of diffusion steps and the placement of the timesteps $({t_{k}})_{k=1}^{n-1}$.
More precisely, Proposition 1 characterizes, at a fixed number of timesteps, the behavior of the reparameterized gradient as ${t_{1}}\to 0$. The proof is given in Appendix A.
Proposition 1.
@@
With $L=1$ and the timesteps $({t_{k}})_{k=2}^{n-1}$ fixed, under assumptions stated in the Appendix, we have for all $\theta\in\Theta$, $\lim_{{t_{1}}\to 0}\big\|\operatorname{J}_{\theta}\!T^{\theta}_{0}(X_{1})\big\|=0\;,\quad\mathbb{P}\text{-a.s.}$ (12) with $X_{1}\sim\mathcal{N}(0,\mathbf{I}_{K})$. The proof consists in showing that, as ${t_{1}}\to 0$, the last DDIM step $T^{\theta}_{0\mathrel{\raisebox{0.45206pt}{$\scriptscriptstyle\mid$}}{t_{1}}}$ collapses almost all points in ${\mathbb{R}}^{K}$ onto a single one-hot vector, and as a consequence, the Jacobian of $T^{\theta}_{0}$ with respect to $\theta$ vanishes. We illustrate Proposition 1 in Figure 1.
@@
Following the previous discussion, $t_{1}$} should not be chosen so small that the gradients become uninformative.
3.3 Extensions
ReinMax extension.